OpenStudy (noahbred):

is the answer 3? 3x – 5 = 7x – 17 A. 6 B. 5 C. 3 D. 2

pooja195 (pooja195):

yep yep :-)

OpenStudy (noahbred):

Simplifying 3x + -5 = 7x + -17 Reorder the terms: -5 + 3x = 7x + -17 Reorder the terms: -5 + 3x = -17 + 7x Solving -5 + 3x = -17 + 7x Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-7x' to each side of the equation. -5 + 3x + -7x = -17 + 7x + -7x Combine like terms: 3x + -7x = -4x -5 + -4x = -17 + 7x + -7x Combine like terms: 7x + -7x = 0 -5 + -4x = -17 + 0 -5 + -4x = -17 Add '5' to each side of the equation. -5 + 5 + -4x = -17 + 5 Combine like terms: -5 + 5 = 0 0 + -4x = -17 + 5 -4x = -17 + 5 Combine like terms: -17 + 5 = -12 -4x = -12 Divide each side by '-4'. x = 3 Simplifying x = 3

pooja195 (pooja195):

nice copy and paste skills

OpenStudy (noahbred):

TY

pooja195 (pooja195):

YW